OpenStudy (anonymous):

write the equation in standard form 6-2(x-y)=3x

OpenStudy (solomonzelman):

6-2(x-y)=3x expand the '-2(x-y)' 6-2x+2y=3x add 2x to both sides 6-2x+2y+2x=3x+2x 6+2y=3x+2x 6+2y=5x subtract 6 from both sides 6+2y-6=5x-6 2y=5x-6 divide each part by 2 2y/2=5x/2-6/2 ---> y=5/2x-3